JDog Corporation owns stock in Oscar Inc. valued at $2,000,000 at the beginning of the year and $2,200,000 at year-end. JDog received a $10,000 dividend from Oscar Inc. What temporary book–tax differences associated with its ownership in Oscar stock will JDog report for the year in the following alternative scenarios (income difference only—ignore the dividends-received deduction)? (Leave no answer blank. Enter zero if applicable.)

a. JDog owns 5 percent of the Oscar Inc. stock. Oscar's income for the year was $500,000.

Book tax difference= _____?

b. JDog owns 40 percent of the Oscar Inc. stock. Oscar's income for the year was $500,000.

Book tax difference=______?
Answer #1

Case a

When the ownership in Oscar Inc. is only 5%, the unrealized gains or losses in value of investments shall be recognized in the books. Such unrealized gain or losses shall not be allowed for tax purposes.

The Increase in value of Oscar Inc. = (2,200,000 - 2,000,000) = $ 200,000
Book tax difference = 200000*5% = $ 10,000

Case b

When the ownership in Oscar Inc. is 40%, equity method shall be followed in recording the investments. The share in net income of Oscar shall be the book tax difference.

Net Income of Oscar Inc. = $ 500,000
Book tax difference = 500000*40% = $ 200,000